# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ 2.3 Properties of Besov–Lorentz spaces
$\leqslant 2^{Nj/p}\sum_{m=1}^{2^{k\log_{2}\Omega}\Omega}\left\|\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k+j})}(y)\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\chi_{B(x+z_{m},1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|$
$\leqslant C2^{k\log_{2}{\Omega}+Nj/p}\left\|\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant C2^{Nj/p}2^{-k}\left\||y|^{1+\log_{2}{\Omega}}\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}\mathrel{}\middle|\mathrel{}L^{\frac{p}{p-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant C2^{Nj/p}2^{-k}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
for each $k\in\mathbb{N}$. Plugging these estimate into (2.3) yields (2.3) for all $j\geqslant 1$ ($j=0$ case also follows by the same argument).
@@
(2) We see $q=\infty$ case (The rest case is proved in a similar way). Clearly
@@
holds for all $\lambda>0$ and for all $f\in L^{\infty}$. Thus, we have
$\left\|\Delta_{j}f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p/\theta,\infty}\right\|$ $=\sup_{\lambda>0}\lambda^{\theta/p}\left[\Delta_{j}f\chi_{B(z,1)}\right]^{\ast}(\lambda)^{\theta}\left[\Delta_{j}f\chi_{B(z,1)}\right]^{\ast}(\lambda)^{1-\theta}$
$\leqslant\left[\sup_{\lambda>0}\lambda^{1/p}\left[\Delta_{j}f\chi_{B(z,1)}\right]^{\ast}(\lambda)\right]^{\theta}\left\|\Delta_{j}f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{\infty}\right\|^{1-\theta}$
$\leqslant\left\|\Delta_{j}f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|^{\theta}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{\infty}\right\|^{1-\theta}$
(2.3) $\leqslant C2^{jN(1-\theta)/p}\left\|\Delta_{j}f\mid L^{p,\infty}_{\mathrm{ul}}\right\|$
for all $z\in\mathbb{R}^{N}$. ∎
Remark 2.11.
Let $1<p<\infty$ and $1\leqslant q\leqslant\infty$. Then the following embeddings
$L^{p,q}_{\mathrm{ul}}\subset L^{p,\infty}_{\mathrm{ul}}\subset M^{p}_{Q}$
hold for all $1\leqslant Q<p$, where $M^{p}_{Q}$ denotes the local Morrey spaces (see Definition 0.1 in [17]). The first one is a standard fact (see Proposition 1.4.10 in [7]). The second one is obtained as follows. For fixed $z\in\mathbb{R}^{N}$ and $R\in\left]0,1\right]$, we find out that
$R^{\frac{N}{p}-\frac{N}{Q}}\left\|f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{Q}\right\|$ $=R^{\frac{N}{p}-\frac{N}{Q}}\left\|f\chi_{B(z,1)}\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{Q}\right\|$
$\leqslant\sup_{R>0}R^{\frac{N}{p}-\frac{N}{Q}}\left\|f\chi_{B(z,1)}\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{Q}\right\|\left(=\left\|f\chi_{B(z,1)}\mathrel{}\middle|\mathrel{}\mathcal{M}^{p}_{Q}\right\|\right)$
[…]
Combining these embeddings with (2.4) in [17] also leads Proposition 2.10.
@@
Fix $1\leqslant p\leqslant\infty$, $0<\kappa<1$ and $r,r_{0},r_{1}\in\left[1,\infty\right]$. Let $s,s_{0},s_{1}$ satisfty $s=(1-\kappa)s_{0}+\kappa s_{1}$. Then we have the formula
$B^{s}_{p,r}=\left(B^{s_{0}}_{p,r_{0}},B^{s_{1}}_{p,r_{1}}\right)_{\kappa,r}.$
In addition, assume $1<p<\infty$. Then we have the formula
$B^{s}_{(p,\infty),r}=\left(B^{s_{0}}_{(p,\infty),r_{0}},B^{s_{1}}_{(p,\infty),r_{1}}\right)_{\kappa,r}.$
@@
See Theorem 6.4.5.(1) in [4]. For the latter, combine Theorem 2.2 in [19] and Definition 4.5 in [19], where we take $E=L^{p,\infty}_{\mathrm{ul}}$. ∎
@@ 2.4 Decay estimates of $S(t)$
In this subsection, we collect the decay estimates of operators $S(t)$ on our function spaces. We begin with Besov–Lorentz spaces.
@@
Let $p\in\left]1,\infty\right[$, $r\in\left[1,\infty\right]$ and $s\geqslant\sigma$. Then, there exists $C>0$ such that the estimate
$\left\|S(t)f\mid B^{s}_{(p,\infty),r}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)\left\|f\mid B^{\sigma}_{(p,\infty),r}\right\|$ (2.5)
holds for all $t>0$ and for all $f\in B^{\sigma}_{(p,\infty),r}$. Furthermore, if $s>\sigma$, the estimate
$\left\|S(t)f\mid B^{s}_{(p,\infty),1}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)\left\|f\mid B^{\sigma}_{(p,\infty),\infty}\right\|$ (2.6)
holds for all $t>0$ and for all $f\in B^{\sigma}_{(p,\infty),\infty}$.
@@
We can prove it in the same way as Theorem 2.1 in [30] with the aid of Proposition 2.3 and 2.12 (see also [1, 18, 17]). ∎
We shall make use of the decay estimates of operators $S(t)$ on Lorentz spaces.
@@
Let $1<p\leqslant q<\infty$. Then, there exists $C>0$ such that the estimate
$\left\|S(t)f\mid L^{q,\infty}_{\mathrm{ul}}\right\|\leqslant C\left(1+t^{\frac{N}{\theta}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)\left\|f\mid L^{p,\infty}_{\mathrm{ul}}\right\|$. (2.7)
holds for all $t>0$ and for all $f\in L^{p,\infty}_{\mathrm{ul}}$.
@@
$p=q$ case. By using Proposition 2.3 and Lemma 2.1 in [21], we gain
$\left\|S(t)f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ $=\left\|\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\right]\ast f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$=\left\|\mathcal{F}^{-1}\left[\exp\left(-|\xi|^{\theta}\right)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|=C\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$.
$p<q$ case. By Propositions 2.9, 2.10 and 2.13, it is seen that
$\left\|S(t)f\mathrel{}\middle|\mathrel{}L^{q,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{0}_{(q,\infty),1}\right\|=C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{\frac{N(q-p)}{qp}-\frac{N(1-p/q)}{p}}_{\left(\frac{p}{p/q},\infty\right),1}\right\|$
$\leqslant C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{\frac{N(q-p)}{qp}}_{(p,\infty),1}\right\|\leqslant C\left(1+t^{\frac{N}{\theta}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)\left\|f\mathrel{}\middle|\mathrel{}B^{0}_{(p,\infty),\infty}\right\|$
$\leqslant C\left(1+t^{\frac{N}{\theta}\left(\frac{1}{q}-\frac{1}{p}\right)}\right)\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
holds for all $t>0$ and for all $f\in L^{p,\infty}_{\mathrm{ul}}$. ∎
@@
One cannot drop “1+1+” in (2.7); one arrives at a contradiction by considering $f(x)\equiv 1$ and taking the limit $t\to\infty$, for instance. This is a “cost” of employing uniformly local function spaces. In this article, however, the term “1+1+” is negligible, as we restrict our attention to local-in-time solvability and take $t<1$.
2.5 An estimate on the forcing term
In this subsection, we establish the following proposition that will be used in the proof of Theorem 1.9; therein lies the novelty of this paper.
@@ 2.5 An estimate on the forcing term
Let $T>0$ and $p\in\left]1,\infty\right[$. There exists a poitive constant $C$ depending on $T$ such that the inequality
[…]
To the best of the author’s knowledge, this estimate seems not to have received attention, even for usual Besov spaces and for $\theta=2$. We note that the proof below does not depend on the base space $L^{p,\infty}_{\mathrm{ul}}$ except for the convolution inequality of Young type (Proposition 2.3).
@@
$\mathcal{F}\left[\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\right](\xi)$ $=\int_{0}^{t}e^{-(t-\tau)|\xi|^{\theta}}\mathcal{F}\mu(\xi)\,d\tau=\frac{1-e^{-|\xi|^{\theta}t}}{|\xi|^{\theta}}\mathcal{F}\mu(\xi).$ (2.9)
@@
(2.9) $=\Phi_{j}(\xi)C_{T}(\xi)\int_{0}^{T}\varphi_{j}(\xi)\mathcal{F}\left[\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\right](\xi)\,dt$
for all $j\in\mathbb{Z}_{\geqslant 1}$. Here $\Phi_{j}$ is the same as in the proof of Proposition 2.10. Thanks to Proposition 2.3, we see that
$\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $=\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\ast\int_{0}^{T}\Delta_{j}\left[\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\right]\,dt\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T}\left\|\Delta_{j}\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\,dt$
@@
$\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T/2^{\theta j}}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T/2^{\theta j}}\left\|\Delta_{j}\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|dt$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T/2^{\theta j}}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|$
$\qquad\qquad\qquad\times\int_{0}^{T/2^{\theta j}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|\,dt.$ (2.10)
Note that we have
@@
$\int_{0}^{\frac{T}{2^{\theta j}}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|dt$ $\leqslant\frac{T}{2^{\theta j}}\sup_{0\leqslant t\leqslant\frac{T}{2^{\theta j}}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|$
$(j\geqslant 0\Rightarrow)\quad\leqslant\,\frac{T}{2^{\theta j}}$ $\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|.$ (2.12)
@@
$2^{-\theta j}\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ $\leqslant C^{\sharp}_{T}\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|$
for all $j\in\mathbb{Z}_{\geqslant 1}$. In a similar way, we can estimate $\left\|\Delta_{0}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|$ as follows:
$\left\|\Delta_{0}\mu\mathrel{}\middle|\mathrel{}L^{p,\infty}_{\mathrm{ul}}\right\|\leqslant T\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|.$
@@
Therefore, taking the supremum of $2^{-\theta j}\left\|\Delta_{j}\mu\mid L^{p,\infty}_{\mathrm{ul}}\right\|$ over $j\in\mathbb{N}$ implies
$\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,\infty),\infty}\right\|$ $\leqslant\left(C^{\flat}_{T}+C^{\sharp}_{T}\right)\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu(x)\,d\tau\mathrel{}\middle|\mathrel{}B^{0}_{\left(p,\infty\right),\infty}\right\|.$

### v2

@@ 2.3 Properties of BLM spaces
$\leqslant 2^{jN/q}\sum_{m=1}^{2^{k\log_{2}\Omega}\Omega}\left\|\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k+j})}(y)\mathrel{}\middle|\mathrel{}L^{\frac{q}{q-1},1}\right\|\left\|\Delta_{j}f\chi_{B(x-z_{m},1)}\mathrel{}\middle|\mathrel{}L^{q,\infty}\right\|$
$\leqslant C2^{k\log_{2}{\Omega}+jN/q}\left\|\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}\mathrel{}\middle|\mathrel{}L^{\frac{q}{q-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant C2^{jN/q}2^{-k}\left\||y|^{1+\log_{2}{\Omega}}\Psi_{0}(y)\chi_{\mathbb{R}^{N}\setminus B(0,2^{k})}\mathrel{}\middle|\mathrel{}L^{\frac{q}{q-1},1}\right\|\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant C2^{jN/q}2^{-k}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
for each $k\in\mathbb{N}$. Plugging these estimate into (2.3) yields (2.3) for all $j\geqslant 1$ ($j=0$ case also follows by the same argument). ∎
Proof of Proposition 2.12.
@@
(2) We see $\kappa=\infty$ case (the rest case is proved in a similar way). Clearly
@@
holds for all $\lambda>0$ and for all $f\in L^{\infty}$. Let $l\in\left]0,1\right[$ and $j\in\mathbb{N}$. Then, we have
$\left\|\Delta_{j}f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{q/l,\infty}\right\|$ $\leqslant\sup_{\lambda>0}\lambda^{l/q}\left[\Delta_{j}f\chi_{B(z,R)}\right]^{\ast}(\lambda)^{l}\left[\Delta_{j}f\chi_{B(z,R)}\right]^{\ast}(\lambda)^{1-l}$
$=\left[\sup_{\lambda>0}\lambda^{1/q}\left[\Delta_{j}f\chi_{B(z,R)}\right]^{\ast}(\lambda)\right]^{l}\left\|\Delta_{j}f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{\infty}\right\|^{1-l}$
$\leqslant\left\|\Delta_{j}f\chi_{B(z,R)}\mathrel{}\middle|\mathrel{}L^{q,\infty}\right\|^{l}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}L^{\infty}\right\|^{1-l}$
(2.3) $\leqslant CR^{\frac{Nl}{q}-\frac{Nl}{p}}2^{jN(1-l)/p}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
for all $z\in\mathbb{R}^{N}$ and $R\in\left]0,1\right]$, and therefore the estimate
$2^{\left(s-\frac{N(1-l)}{p}\right)j}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p/l}_{q/l,\infty}\right\|\leqslant C2^{sj}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
holds for all $s\in\mathbb{R}$ and $j\in\mathbb{N}$. ∎
@@
Fix $1\leqslant p\leqslant\infty$, $0<\kappa<1$ and $r,r_{0},r_{1}\in\left[1,\infty\right]$. Let $s,s_{0},s_{1}$ satisfy $s=(1-\kappa)s_{0}+\kappa s_{1}$. Then we have the formula
$B^{s,r}_{p}=\left(B^{s_{0},r_{0}}_{p},B^{s_{1},r_{1}}_{p}\right)_{\kappa,r}.$
@@
See Theorem 6.4.5.(1) in [4]. ∎
@@ 2.4 Decay estimates of $S(t)$
In this subsection, we collect the decay estimates of operators $S(t)$ on our function spaces. We begin with BLM spaces.
@@
Let $1<q\leqslant p<\infty$, $r\in\left[1,\infty\right]$ and $s\geqslant\sigma$. Then there exists $C>0$ such that the estimate
$\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{s,r}_{p,q,\infty}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)\left\|f\mathrel{}\middle|\mathrel{}B^{\sigma,r}_{p,q,\infty}\right\|$ (2.5)
holds for all $t>0$ and for all $f\in B^{\sigma,r}_{p,q,\infty}$. Furthermore, if $s>\sigma$, the estimate
$\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{s,1}_{p,q,\infty}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)\left\|f\mathrel{}\middle|\mathrel{}B^{\sigma,\infty}_{p,q,\infty}\right\|$ (2.6)
holds for all $t>0$ and for all $f\in B^{\sigma,\infty}_{p,q,\infty}$.
@@
First we prove (2.5). It suffices to show that the following inequality
$\left\|\Delta_{j}S(t)f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant C\left(1+t^{\frac{-(s-\sigma)}{\theta}}\right)2^{-(s-\sigma)j}\left\|\Delta_{j}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
holds for all $j\in\mathbb{N}$. The case $j\geqslant 1$ can be treated in the same way as in the proof of Theorem 2.9 in [20], relying on Proposition 2.3 (see also [1, 21, 33]). For $j=0$, the inequality follows by combining Proposition 2.3 with Lemma 2.1 in [25], as shown below:
$\left\|\Delta_{0}S(t)f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $=\left\|\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\right]\ast\Delta_{0}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\exp\left(-t|\xi|^{\theta}\right)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\left\|\Delta_{0}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|=C\left\|\Delta_{0}f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|.$
To obtain (2.6), we fix arbitrary $a>s$ and $m\in\mathbb{N}$, and perform the following calculation with (2.5):
$\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{s,1}_{p,q,\infty}\right\|$ $\leqslant\sum_{j=0}^{m-1}2^{(s-\sigma)j}\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{\sigma,\infty}_{p,q,\infty}\right\|+\sum_{j=m}^{\infty}2^{-(a-s)j}\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{a,\infty}_{p,q,\infty}\right\|$
$\leqslant C\left(2^{(s-\sigma)m}+2^{-(a-s)m}\left(1+t^{-\frac{a-\sigma}{\theta}}\right)\right)\left\|f\mathrel{}\middle|\mathrel{}B^{\sigma,\infty}_{p,q,\infty}\right\|.$
Optimizing in $m$, we get the claim. ∎
We confirm the decay estimates of operators $S(t)$ on Lorentz–Morrey spaces.
@@
Let $1<q\leqslant p<\infty$ and $l\in\left]0,1\right]$. Then there exists $C>0$ such that the estimate
$\left\|S(t)f\mathrel{}\middle|\mathrel{}M^{p/l}_{q/l,\infty}\right\|\leqslant C\left(1+t^{-\frac{N(1-l)}{\theta p}}\right)\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|.$ (2.7)
holds for all $t>0$ and for all $f\in M^{p}_{q,\infty}$.
@@
Case $l<1$. By Propositions 2.11, 2.12 and 2.15, it is seen that
$\left\|S(t)f\mathrel{}\middle|\mathrel{}M^{p/l}_{q/l,\infty}\right\|$ $\leqslant C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{0,1}_{p/l,q/l,\infty}\right\|\leqslant C\left\|S(t)f\mathrel{}\middle|\mathrel{}B^{N(1-l)/p,1}_{p,q,\infty}\right\|$
$\leqslant C\left(1+t^{-\frac{N(1-l)}{\theta p}}\right)\left\|f\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|\leqslant C\left(1+t^{-\frac{N(1-l)}{\theta p}}\right)\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
holds for all $t>0$ and for all $f\in M^{p}_{q,\infty}$.
Case $l=1$. This is done in the proof of the estimate (2.5) ($j=0$ case). ∎
@@
One cannot drop “1+1+” in (2.7); one arrives at a contradiction by considering $f(x)\equiv 1$ and taking the limit $t\to\infty$, for instance. In this article, however, the term “1+1+” is negligible, as we restrict our attention to local-in-time solvability and make use of the following inequality:
$\left(1+t^{-\frac{N(1-l)}{\theta p}}\right)\chi_{\left]0,T\right[}(t)\leqslant Ct^{-\frac{N(1-l)}{\theta p}}\quad\text{for }0<T\leqslant 1\,\text{ and for }t>0.$
2.5 An estimate on the inhomogeneous term
In this subsection, we establish the following proposition that will be used in the proof of Theorem 1.10; this constitutes one of the novelties of this paper.
@@ 2.5 An estimate on the inhomogeneous term
Let $T>0$ and $1<q\leqslant p<\infty$. There exists a positive constant $C$ depending on $T$ such that the inequality
$\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta,\infty}_{p,q,\infty}\right\|\leqslant C\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}L^{\infty}(0,T;M^{p}_{q,\infty})\right\|$
holds for all $\mu\in\left\{f\in\mathcal{S}^{\prime}(\mathbb{R}^{N})\mathrel{}\middle|\mathrel{}\int_{0}^{t}S(t-\tau)f\,d\tau\in L^{\infty}(0,T;M^{p}_{q,\infty})\right\}$.
To the best of the author’s knowledge, this estimate seems not to have received attention, even for usual Besov spaces and for $\theta=2$. We note that the proof below does not depend on the base space $M^{p}_{q,\infty}$ except for the convolution inequality of Young type (Proposition 2.3).
@@
$\mathcal{F}\left[\int_{0}^{t}S(t-\tau)\mu\,d\tau\right](\xi)$ $=\int_{0}^{t}e^{-(t-\tau)|\xi|^{\theta}}\mathcal{F}\mu(\xi)\,d\tau=\frac{1-e^{-|\xi|^{\theta}t}}{|\xi|^{\theta}}\mathcal{F}\mu(\xi).$ (2.9)
@@
(2.9) $=\Phi_{j}(\xi)C_{T}(\xi)\int_{0}^{T}\varphi_{j}(\xi)\mathcal{F}\left[\int_{0}^{t}S(t-\tau)\mu\,d\tau\right](\xi)\,dt$
for all $j\in\mathbb{Z}_{\geqslant 1}$. Here $\Phi_{j}$ is the same as in the proof of Lemma 2.13. Thanks to Proposition 2.3, we see that
$\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $=\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\ast\int_{0}^{T}\Delta_{j}\left[\int_{0}^{t}S(t-\tau)\mu\,d\tau\right]\,dt\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T}\left\|\Delta_{j}\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\,dt$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|\,dt$
@@
$\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$
$\leqslant\left\|\mathcal{F}^{-1}\left[\Phi_{j}(\xi)C_{T/2^{\theta j}}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\int_{0}^{T/2^{\theta j}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|\,dt.$ (2.10)
Note that
@@
$\int_{0}^{\frac{T}{2^{\theta j}}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|dt$ $\leqslant\frac{T}{2^{\theta j}}\sup_{0\leqslant t\leqslant\frac{T}{2^{\theta j}}}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|$
$\leqslant\,\frac{T}{2^{\theta j}}$ $\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|.$ (2.12)
@@
$2^{-\theta j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ $\leqslant C^{\sharp}_{T}\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|$
for all $j\in\mathbb{Z}_{\geqslant 1}$. In a similar way, we can estimate $\left\|\Delta_{0}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ as follows:
$\left\|\Delta_{0}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant T\left\|\mathcal{F}^{-1}\left[\Phi_{(0)}(\xi)C_{T}(\xi)\right]\mathrel{}\middle|\mathrel{}L^{1}\right\|\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|.$
@@
Therefore, taking the supremum of $2^{-\theta j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|$ over $j\in\mathbb{N}$ implies
$\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta,\infty}_{p,q,\infty}\right\|$ $\leqslant\left(C^{\flat}_{T}+C^{\sharp}_{T}\right)\sup_{0\leqslant t\leqslant T}\left\|\int_{0}^{t}S(t-\tau)\mu\,d\tau\mathrel{}\middle|\mathrel{}B^{0,\infty}_{p,q,\infty}\right\|.$